# Key Words.: stars: abundances – Local Group – galaxies: dwarf – galaxies: formation1 Introduction

### v1

@@ 3.3 Final stellar parameters and abundance determination
We determined the stellar chemical abundances via the measurement of the equivalent widths (EWs) or spectral synthesis of atomic transition lines, when necessary, for example in the case of blended lines. Lines present in the spectra and our line list are detected and their EWs measured with DAOSPEC (Stetson & Pancino, 2008). This code performs a Gaussian fit of each individual line and measures its corresponding EW. Although DAOSPEC fits saturated Gaussians to strong lines, it cannot fit the wider Lorentz-like wings of the profile of very strong lines, in particular beyond 120 mÅ at very high resolution (Kirby & Cohen, 2012). For some of the strongest lines in our spectra, we therefore derived the abundances by spectral synthesis (see below).
@@
The measured EWs are provided in Table 7. Values in bracket indicate that the corresponding abundances were derived by spectral synthesis. The abundance derivation from EWs and the spectral synthesis calculation were performed with the Turbospectrum code (Alvarez & Plez, 1998; Plez, 2012), which assumes local thermodynamic equilibrium (LTE), but treats continuum scattering in the source function. We used a plane-parallel transfer for the line computation; this is consistent with our previous work on EMP stars (Tafelmeyer et al., 2010; Jablonka et al., 2015; Lucchesi et al., 2020).
@@
In order to derive the final T and the microturbulence velocities (), we checked or required no trend between the abundances derived from Fe i and excitation potential () or the predicted333According to Magain (1984), the use of observed EWs produce an increase of by 0.10.2 , which would be reflected in a decrease of the measured [Fe/H] values by a few hundredths of a dex in a systematic way. A variation like this does not change the results in a significant way. EWs. In order to minimize the NLTE effect on the measured abundances we excluded from the analysis Fe i lines with ¡ 1.4 eV. Furthermore very strong lines (EW ¿ 120mÅ) with strong wings that cannot be well fitted are also excluded.
@@
Starting from the initial photometric parameters of Table 3, we adjusted and by minimizing the slopes of the diagnostic plots, within its 2 uncertainty. We did not force ionization equilibrium between Fe I and Fe II, taking into account that there will likely be NLTE effects at these low metallicities (Amarsi et al., 2016; Mashonkina et al., 2017a; Ezzeddine et al., 2017). For each iteration the corresponding values of were computed from its relation with (Eq. 1), assuming the updated values of , and adjusting the model metallicity to the mean iron abundance derived in the previous iteration. The final values of T are less than 30 K away from the initial photometric estimates, at the exception of fnx0579x1 which is 138 K cooler than the mean photometric temperature.
@@
We derived the chemical abundances of the strong lines with measured EW ¿ 100 mÅ by spectral synthesis. These abundances were obtained using our own code (as in Lucchesi et al. 2020, 2022), which performs a -minimization between the observed spectral features and a grid of synthetic spectra calculated on the fly with Turbospectrum. A line of a chemical element is synthesized in a wavelength range of 50 Å. It is optimized by varying its abundance in steps of 0.1 dex, from [/Fe] = dex to [/Fe] = dex. In the same way, the resolution of the synthetic spectra is optimized when needed. Starting from the nominal instrumental resolution, synthetic spectra can be convolved with a wide range of Gaussian widths for each abundance step. A second optimization, with abundance steps of 0.01 dex, is then performed in a smaller range around the minimum in order to refine the results. Similarly, the elements with a significant hyperfine structure (HFS; e.g. Sc, Mn, Co, and Ba) have been determined by running Turbospectrum in its spectral synthesis mode in order to properly take into account blends and the HFS components in the abundance derivation, as in North et al. (2012), Prochaska & McWilliam (2000) for Sc and Mn, and from the Kurucz web site444http://kurucz.harvard.edu/linelists.html for Co and Ba.
The final abundances are listed in Table 4. The solar abundances are taken from Asplund et al. (2009).
Fe I | Fe II | C** | O I | Na I | Mg I | Al I | Si I | Ca I | Sc II | Ti I | Ti II | Cr I | Mn I | Co I | Ni I | Cu I | Zn I | Sr II | Y II | Zr II | Ba II | La II | Pr II | Nd II | Eu II | Dy II | |
log(X) | 7.50 | 7.50 | 8.43 | 8.69 | 6.24 | 7.60 | 6.45 | 7.51 | 6.34 | 3.15 | 4.95 | 4.95 | 5.64 | 5.43 | 4.99 | 6.22 | 4.19 | 4.56 | 2.87 | 2.21 | 2.58 | 2.18 | 1.10 | 0.72 | 1.42 | 0.52 | 1.10 |
fnx_06_019 | |||||||||||||||||||||||||||
No. lines* | 67 | 2 | 1 | 1 | 2 | 3 | 1 | 1 | 6 | 6 | 9 | 7 | 4 | 3 | 1 | 3 | 1 | 1 | 1 | 5 | 1 | 4 | 4 | 1 | 3 | 2 | 3 |
log(X) | ¡ | ||||||||||||||||||||||||||
[X/H] | ¡ | ||||||||||||||||||||||||||
[X/Fe] | ¡ | ||||||||||||||||||||||||||
Error | |||||||||||||||||||||||||||
fnx0579x–1 | |||||||||||||||||||||||||||
No. lines* | 76 | 4 | 1 | 1 | 2 | 4 | 1 | 11 | 5 | 10 | 8 | 5 | 3 | 1 | 4 | 1 | 1 | 2 | 4 | ||||||||
log(X) | |||||||||||||||||||||||||||
[X/H] | |||||||||||||||||||||||||||
[X/Fe] | |||||||||||||||||||||||||||
Error | |||||||||||||||||||||||||||
car1_t174 | |||||||||||||||||||||||||||
No. lines* | 10 | 2 | 1 | 1 | 2 | 1 | 1 | 1 | 1 | 2 | |||||||||||||||||
log(X) | |||||||||||||||||||||||||||
[X/H] | |||||||||||||||||||||||||||
[X/Fe] | |||||||||||||||||||||||||||
Error | |||||||||||||||||||||||||||
car1_t194 | |||||||||||||||||||||||||||
No. lines* | 24 | 2 | 1 | 2 | 2 | 2 | 3 | 3 | 1 | 1 | 1 | 2 | |||||||||||||||
log(X) | |||||||||||||||||||||||||||
[X/H] | |||||||||||||||||||||||||||
[X/Fe] | |||||||||||||||||||||||||||
Error | |||||||||||||||||||||||||||
car1_t200 | |||||||||||||||||||||||||||
No. lines* | 12 | 1 | 2 | 2 | 2 | 1 | 1 | 1 | 2 | ||||||||||||||||||
log(X) | |||||||||||||||||||||||||||
[X/H] | |||||||||||||||||||||||||||
[X/Fe] | |||||||||||||||||||||||||||
Error | |||||||||||||||||||||||||||
LG04c_0008 | |||||||||||||||||||||||||||
No. lines* | 35 | 6 | 1 | 2 | 2 | 1 | 2 | 1 | 1 | 2 | 1 | 1 | 2 | ||||||||||||||
log(X) | |||||||||||||||||||||||||||
[X/H] | |||||||||||||||||||||||||||
[X/Fe] | |||||||||||||||||||||||||||
Error |
@@ 3.4 Error budget
The uncertainties on the abundances were derived considering the uncertainties on the atmospheric parameters and on the EWs, in a similar procedure to our other works (e.g. Tafelmeyer et al., 2010; Jablonka et al., 2015; Hill et al., 2019; Lucchesi et al., 2020).
1.
Uncertainties due to the atmospheric parameters. To estimate the sensitivity of the derived abundances to the adopted atmospheric parameters, we repeated the abundance analysis and varied only one stellar atmospheric parameter at a time by its corresponding uncertainty, keeping the others fixed and repeating the analysis. The estimated internal errors are 100 K in T, 0.15 dex in (g), and 0.15 km s in for the UVES sample, and 150 K in T, 0.15 dex in (g) for the X-Shooter sample. Because the S/N and atmospheric parameters of our sample stars are very close to each other, we estimated the typical errors considering a single reference star. Table 5 list the effects of these changes on the derived abundances for fnx_06_019 in the UVES sample, and Table 6 list the effects of these changes on the derived abundances for carLG04c_0008 in the X-Shooter sample.
2.
Uncertainties due to EWs or spectral fitting. The uncertainties on the individual EW measurements are provided by DAOSPEC (see Table 7) and computed according to the following formula (Stetson & Pancino, 2008) :
(2) where and are the intensity of the observed line profile at pixel and its uncertainty, and and are the intensity and uncertainty of the corresponding continuum. The uncertainties on the intensities are estimated from the scatter of the residuals that remain after subtraction of the fitted line (or lines, in the case of blends). The corresponding uncertainties on individual line abundances are propagated by Turbospectrum. This is a lower limit to the real EW error because systematic errors like the continuum placement are not accounted for.
In order to account for additional sources of error, we quadratically added a 5% error to the EW uncertainty, so that no EW has an error smaller than 5%. For the abundances derived by spectral synthesis (e.g., strong lines, hyperfine structure, or carbon from the G band), the uncertainties were visually estimated by gradually changing the parameters of the synthesis until the deviation from the observed line became noticeable.
The abundance uncertainty for an element X due to the individual EW uncertainties ( propagated from ) are computed as :
where represents the number of lines measured for element X.
The dispersion around the mean abundance of an element X measured from several lines is computed as :
@@
Ba II |
3.5 Specific comments on the abundance determination
@@ 3.5.2 Lithium
Unfortunately the Li resonance doublet at 6707 Å was not measurable in any of our stars at our available quality of the UVES or X-Shooter spectra. This is expected, since Li on the surface of RGB stars is typically depleted due to dredge-up and mixing (e.g. Gratton et al., 2000; Lind et al., 2009). None of our target stars are therefore lithium-enhanced giants, which have been found in small fractions in most types of environments, including dwarf galaxies (e.g. Kirby et al., 2012; Hill et al., 2019).
@@ 3.5.3 Sodium and Aluminium
The sodium was measured from the two 5889.951 Å and 5895.924 Å resonance lines through spectral synthesis.
The X-Shooter spectra of the four Carina target stars are presented in the region of the Na lines and the CH absorption in Fig. 4 and Fig. 4 respectively. They are compared to a star from Starkenburg et al. (2013) in the Sculptor dSph galaxy, Scl074_02, with very similar atmospheric parameters (T=4595 K g=1.21 [Fe/H]=3.0 =2.1 km s), but significantly lower Na. The carbon abundance is very similar in all stars shown in Fig. 4, with low [C/Fe] . However, the difference can be clearly seen in the sodium region (Fig. 4), where the sodium lines appear stronger in Carina than in the Sculptor star. We note that interstellar Na lines are observed at different radial velocities and are unlikely contaminating the measured lines.
@@
. The UVES Mg abundances are based on three lines. Two of them are rather strong (5172.684 and 5183.604 Å), with EW ¿ 100 mÅ and have non-Gaussian line profiles. The abundances of these lines are not consistent with the weaker 5528.405 Å line. For this reason, we derived the Mg abundance through spectral synthesis, after which all lines had consistent abundances. Three additional Mg i lines (4167.271, 4351.906, and 5711.088 Å) are detected in our spectra, but were discarded because they are too noisy, strongly blended and too weak, respectively. The Mg abundances from the X-Shooter sample are obtained through spectral synthesis in two 20 Å windows, one centered on the 5172.684 Å line, taking into account the blends at the X-Shooter resolution, and one centered on the 5183.604 Å line.
. One Si I line is detected in our UVES spectra, at 4102.936 Å, but in a noisy part of the spectrum and it falls closely to the strong H absorption line. However, we were able to derive a Si abundance in both Fornax stars, based on this one line.
. Three Ca I lines were used in the X-Shooter spectra for the Carina stars. Calcium was not measurable in the star car1_t200. In the Fornax sample, in the red part of the UVES spectra ( ¿5500 Å ) six and eleven clean Ca I lines were used, for fnx0579x-1 and fnx_06_019, respectively. Examples of the Ca I lines for fnx_06_019 are shown in Fig. 5.
@@ 3.5.5 Iron-peak elements
. Cr abundances are derived from 4 to 5 Cr I lines in the red part of the UVES spectra, all of them give consistent results from their EW. Five extra lines are detected in our spectra (4254.352, 4274.812, 4289.73, 5206.023 and 5208.409 Å) but they were rejected for being too strong (¿110 mÅ) or too noisy. In the X-Shooter spectra only the strongest 5206.023 and 5208.409 Å lines were accessible, Cr abundances were obtained from a single spectral synthesis in a 20 Å wavelength range covering the two lines and taking into account the blends at this resolution.
. Mn abundances rely on the three Mn I 4030.75 Å 4033.06 Å and 4034.48 Å lines. These were synthesized taking into account their HFS components and give consistent abundance results. The Mn lines 4041.35 Å and 4823.52 Å are also present in our spectra but they are weak (30 mÅ) and too much affected by the noise and were thus discarded. In the case of X-Shooter spectra, a single spectral synthesis has been done in a 20 Å window centered on the Mn triplet.
@@ 3.5.6 Neutron-capture elements
. Three Sr II lines are present in our spectra (4077.709, 4161.792 and 4215.519 Å). While the 4215.519 Å line was the most reliable in the UVES spectra, the 4077.709 Å line was less affected by blends and better suited for the X-Shooter spectra.
. Ba abundances were measured from 2 to 4 lines (4554.029, 4934.076, 5853.668, 6141.713 and 6496.897 Å) in the UVES and X-Shooter spectra by spectral synthesis taking into account blends and HFS.
Additional neutron-capture elements were measurable in the spectrum of the -process rich star fnx_06_019 in Fornax. Fig. 6 compares two UVES spectra in the region of lanthanum and europium lines, showing a clear neutron-capture enhancement in fnx_06_019 (see also Fig. 15).
. The La abundance was determined from the 4920.98 Å line which is the reddest available La II line and the least affected by noise. The detection was checked by computing synthetic spectra for the 4077.34, 4086.71, and 4123.23 Å lines using the abundance derived from the 4920.98 Å line (Fig. 15). All lines were in good agreement with the adopted abundance.
@@ 4.1 Literature comparison samples
In Figures 7–13, abundances from the Milky Way are shown with gray dots and are compiled from Placco et al. (2014) for carbon, from Bensby et al. (2014); Yong et al. (2013) for Na, Mg, Ca, Ti, with the addition of Frebel et al. (2010) for Na. Milky Way iron-peak elements (Sc, Cr, Mn, Co, Ni, Zn) are from Bensby et al. (2014); Yong et al. (2013); Frebel et al. (2010), while abundances of the neutron-capture elements strontium and barium are taken from Roederer (2013). Milky Way europium abundances are from Frebel et al. (2010); Venn et al. (2004).
Abundances measured in Carina are taken from Koch et al. (2008), Venn et al. (2012), and Norris et al. (2017). Norris et al. (2017) compiled 63 RGB stars, from which 14 stars were totally new observations, while 18 stars were re-observations and 31 stars were re-analysis of stars initially studied by Lemasle et al. (2012); Shetrone et al. (2003) and Venn et al. (2012). For the purpose of homogeneity we show the Norris et al. (2017) results and refer to this paper for a detailed description of the individual samples. The only exception is made for the Venn et al. (2012) sample. Since their work includes more chemical elements than Norris et al. (2017), we keep the original results of the nine Carina members from Venn et al. (2012). Fornax comparison samples are taken from Letarte et al. (2006, 2010); Lemasle et al. (2014); Tafelmeyer et al. (2010).

### v2

@@ 3.3 Final stellar parameters and abundance determinations
We determined the stellar chemical abundances via the measurement of the equivalent widths (EWs) or spectral synthesis of atomic transition lines when necessary (see below). Lines present in the spectra and in our line list were detected and their EWs measured with DAOSPEC (Stetson & Pancino, 2008). This code performs a Gaussian fit of each individual line and measures its corresponding EW. Although DAOSPEC fits saturated Gaussians to strong lines, it cannot fit the wider Lorentz-like wings of the profile of very strong lines, in particular beyond 120 mÅ at very high resolution (Kirby & Cohen, 2012). For some of the strongest lines in our spectra, we therefore derived the abundances by spectral synthesis. All other detected lines by DAOSPEC were also inspected visually and the measured EWs were rejected when too uncertain (for example in the case of upper limits the abundances were determined through spectral synthesis), or the lines were totally rejected when the continuum level was too uncertain.
@@
The measured EWs are provided in Table 6 and 3. The values in parentheses indicate that the corresponding abundances were derived by spectral synthesis. The abundance derivation from EWs and the spectral synthesis calculation were performed with the Turbospectrum code (Alvarez & Plez, 1998; Plez, 2012), which assumes local thermodynamic equilibrium (LTE), but treats continuum scattering in the source function. We used a plane-parallel transfer for the line computation; this is consistent with our previous work on EMP stars (Tafelmeyer et al., 2010; Jablonka et al., 2015; Lucchesi et al., 2020).
@@
In order to derive the final Teff and the microturbulence velocities (), we checked or required no trend between the abundances derived from Fe i and excitation potential () or the predicted333According to Magain (1984), the use of observed EWs produce an increase in of 0.10.2 , which would be reflected in a decrease in the measured [Fe/H] values of a few hundredths of a dex in a systematic way. A variation like this does not change the results significantly. EWs. In order to minimize the NLTE effect on the measured abundances we excluded from the analysis Fe i lines with ¡ 1.4 eV. Furthermore, very strong lines (EW ¿ 120mÅ) with strong wings that cannot be well fit are also excluded.
@@
Starting from the initial photometric parameters of Table 3, we adjusted and by minimizing the slopes of the diagnostic plots, within its 2 uncertainty. We did not force ionization equilibrium between Fe I and Fe II, taking into account that there will likely be NLTE effects at these low metallicities (Amarsi et al., 2016; Mashonkina et al., 2017a; Ezzeddine et al., 2017). For each iteration the corresponding values of were computed from its relation with (Eq. 1), assuming the updated values of , and adjusting the model metallicity to the mean iron abundance derived in the previous iteration. The final values of Teff are less than 30 K away from the initial photometric estimates; the only exception is fnx0579x1, which is 138 K cooler than the mean photometric temperature.
@@
We derived the chemical abundances of blended lines, strong lines (EW ¿ 100 mÅ), or upper limits (EW ¡ 25 mÅ) by spectral synthesis. These abundances were obtained using our own code (as in Lucchesi et al. 2020, 2022), which performs a -minimization between the observed spectral features and a grid of synthetic spectra calculated on the fly with Turbospectrum. A line of a chemical element is synthesized in a wavelength range of 10 to 50 Å. It is optimized by varying its abundance in steps of 0.1 dex, from [/Fe] = dex to [/Fe] = dex. In the same way, the resolution of the synthetic spectra is optimized when needed. Starting from the nominal instrumental resolution, synthetic spectra can be convolved with a wide range of Gaussian widths for each abundance step. A second optimization, with abundance steps of 0.01 dex, is then performed in a smaller range around the minimum in order to refine the results. Similarly, the elements with a significant hyperfine structure (HFS; e.g., Sc, Mn, Co, and Ba) are determined by running Turbospectrum in its spectral synthesis mode in order to properly take into account blends and the HFS components in the abundance derivation, as in North et al. (2012) and Prochaska & McWilliam (2000) for Sc and Mn, and from the Kurucz web site444http://kurucz.harvard.edu/linelists.html for Co and Ba.
@@ 3.4 Error budget
The uncertainties on the abundances were derived considering the uncertainties on the atmospheric parameters and on the EWs, in a procedure similar to that used in our other works (e.g., Tafelmeyer et al., 2010; Jablonka et al., 2015; Hill et al., 2019; Lucchesi et al., 2020).
Uncertainties due to the atmospheric parameters. To estimate the sensitivity of the derived abundances to the adopted atmospheric parameters, we repeated the abundance analysis and varied only one stellar atmospheric parameter at a time by its corresponding uncertainty, keeping the others fixed and repeating the analysis. The estimated internal errors are 100 K in Teff, 0.15 dex in (g), and 0.15 km s-1 in for the UVES sample, and 150 K in Teff, 0.15 dex in (g) for the X-Shooter sample. Because the S/N and atmospheric parameters of our sample stars are very close to each other, we estimated the typical errors considering a single reference star. Table 4 lists the effects of these changes on the derived abundances for fnx_06_019 in the UVES sample, and Table 5 lists the effects of these changes on the derived abundances for carLG04c_0008 in the X-Shooter sample.
Uncertainties due to EWs or spectral fitting. The uncertainties on the individual EW measurements () are provided by DAOSPEC (see Table 6) and are computed according to the formula (Stetson & Pancino, 2008)
where and are respectively the intensity of the observed line profile at pixel and its uncertainty, and and are the intensity and uncertainty of the corresponding continuum. The uncertainties on the intensities are estimated from the scatter of the residuals that remain after subtraction of the fitted line (or lines in the case of blends). The corresponding uncertainties on individual line abundances are propagated by Turbospectrum. This is a lower limit to the real EW error because systematic errors, such as the continuum placement, are not accounted for.
In order to account for additional sources of error, we quadratically added a 5% error to the EW uncertainty so that no EW has an error smaller than 5%. For the abundances derived by spectral synthesis (e.g., strong lines, hyperfine structure, or carbon from the G band), the uncertainties were visually estimated by gradually changing the parameters of the synthesis until the deviation from the observed line became noticeable. The abundance uncertainty for an element X due to the individual EW uncertainties ( propagated from ) are computed as
where represents the number of lines measured for element X. The dispersion around the mean abundance of an element X measured from several lines is computed as
@@
| Ba II | ||||
3.5 Specific comments on abundance determinations
@@ 3.5.2 Lithium
Unfortunately, the Li resonance doublet at 6707 Å was not measurable in any of our stars at our available quality of the UVES or X-Shooter spectra. This was expected, since Li on the surface of RGB stars is typically depleted due to dredge-up and mixing (e.g., Gratton et al., 2000; Lind et al., 2009). None of our target stars are therefore lithium-enhanced giants, which have been found in small fractions in most types of environments, including dwarf galaxies (e.g., Kirby et al., 2012; Hill et al., 2019).
@@ 3.5.3 Sodium and aluminium
The sodium was measured from the two 5889.951 Å and 5895.924 Å resonance lines through spectral synthesis. The X-Shooter spectra of the four Carina target stars are presented in the region of the Na lines and the CH absorption in Fig. 4 and Fig. 4, respectively. They are compared to a star from Starkenburg et al. (2013) in the Sculptor dSph galaxy, Scl074_02, with very similar atmospheric parameters (Teff=4595 K g=1.21 [Fe/H]=3.0 =2.1 km s-1), but significantly lower Na. The carbon abundance is very similar in all stars shown in Fig. 4, with low [C/Fe] . However, the difference can be clearly seen in the sodium region (Fig. 4), where the sodium lines appear stronger in Carina than in Sculptor. We note that interstellar Na lines are observed at different radial velocities and are unlikely to be contaminating the measured lines.
@@
. The UVES Mg abundances are based on three lines. Two of them are rather strong (5172.684 and 5183.604 Å), with EW ¿ 100 mÅ and have non-Gaussian line profiles. The abundances of these lines are not consistent with the weaker 5528.405 Å line. For this reason, we derived the Mg abundance through spectral synthesis, after which all lines had consistent abundances. Three additional Mg i lines (4167.271, 4351.906, and 5711.088 Å) were detected in our spectra, but were discarded because too noisy, strongly blended, and too weak, respectively. The Mg abundances from the X-Shooter sample were obtained through spectral synthesis in two 20 Å windows, one centered on the 5172.684 Å line, taking into account the blends at the X-Shooter resolution, and one centered on the 5183.604 Å line.
. One Si I line was detected in our UVES spectra, at 4102.936 Å, but in a noisy part of the spectrum; it falls close to the strong H absorption line. However, we were able to derive a Si abundance in both Fornax stars based on this one line.
. Three Ca I lines were used in the X-Shooter spectra for the Carina stars. Calcium was not measurable in the star car1_t200. In the Fornax sample, in the red part of the UVES spectra ( ¿5500 Å ) 6 and 11 clean Ca I lines were used, for fnx0579x-1 and fnx_06_019, respectively. Examples of the Ca I lines for fnx_06_019 are shown in Fig. 5.
@@ 3.5.5 Iron-peak elements
. Cr abundances were derived from 4 to 5 Cr I lines in the red part of the UVES spectra; all of them give consistent results from their EWs. Five extra lines were detected in our spectra (4254.352, 4274.812, 4289.73, 5206.023, and 5208.409 Å), but they were rejected for being too strong (¿110 mÅ) or too noisy. In the X-Shooter spectra only the strongest 5206.023 and 5208.409 Å lines were accessible. Cr abundances were obtained from a single spectral synthesis in a 20 Å wavelength range covering the two lines and taking into account the blends at this resolution.
. Mn abundances rely on the three Mn I 4030.75 Å, 4033.06 Å, and 4034.48 Å lines. These were synthesized taking into account their HFS components and give consistent abundance results. The Mn lines 4041.35 Å and 4823.52 Å are also present in our spectra, but they are weak (30 mÅ) and too affected by noise, and were thus discarded. In the case of X-Shooter spectra, a single spectral synthesis was done in a 20 Å window centered on the Mn triplet.
@@ 3.5.6 Neutron-capture elements
. Three Sr II lines are present in our spectra (4077.709, 4161.792, and 4215.519 Å). While the 4215.519 Å line was the most reliable in the UVES spectra, the 4077.709 Å line was less affected by blends and better suited for the X-Shooter spectra.
. Ba abundances were measured from two to four lines (4554.029, 4934.076, 5853.668, 6141.713, and 6496.897 Å) in the UVES and X-Shooter spectra by spectral synthesis taking into account blends and HFS.
Additional neutron-capture elements were measurable in the spectrum of the -process rich star fnx_06_019 in Fornax. Figure 6 compares two UVES spectra in the region of the lanthanum and europium lines, showing a clear neutron-capture enhancement in fnx_06_019 (see also Fig. 15).
. The La abundance was determined from the 4920.98 Å line, which is the reddest available La II line and the least affected by noise.
The detection was checked by computing synthetic spectra for the 4077.34, 4086.71, and 4123.23 Å lines using the abundance derived from the 4920.98 Å line (Fig. 15). All the lines were in good agreement with the adopted abundance.
@@ 4.1 Literature comparison samples
In Figures 7–13, the abundances from the Milky Way are shown as gray dots and are compiled from Placco et al. (2014) for carbon; from Bensby et al. (2014), and Yong et al. (2013) for Na, Mg, Ca, Ti; and from Frebel et al. (2010) for Na. The Milky Way iron-peak elements (Sc, Cr, Mn, Co, Ni, Zn) are from Bensby et al. (2014), Yong et al. (2013), and Frebel et al. (2010); while the abundances of the neutron-capture elements strontium and barium are from Roederer (2013). The Milky Way europium abundances are from Frebel et al. (2010), and Venn et al. (2004).
Abundances measured in Carina are from Koch et al. (2008), Venn et al. (2012), and Norris et al. (2017). Norris et al. (2017) compiled 63 RGB stars, of which 14 stars were totally new observations, while 18 stars were re-observations and 31 stars were re-analyses of stars initially studied by Lemasle et al. (2012), Shetrone et al. (2003), and Venn et al. (2012). For the purpose of homogeneity we show the Norris et al. (2017) results and refer to this paper for a detailed description of the individual samples. The only exception is made for the Venn et al. (2012) sample. Since their work includes more chemical elements than Norris et al. (2017), we kept the original results of the nine Carina members from Venn et al. (2012). The Fornax comparison samples are taken from Letarte et al. (2006), Letarte et al. (2010), Lemasle et al. (2014), and Tafelmeyer et al. (2010).